# Hatching-Box: Monitoring the Rearing Process of Drosophila Using an Embedded Imaging and in-vial Detection System

### v1

@@ 4.2.4 Tracking
To analyse the images recorded by the HB-client we implemented an adapted tracking algorithm, which is accessible in the HB-server. This algorithm firstly separates the image into individual rearing vial crops using a one-time calibration step, which automatically generates a mask for each vial (see Appendix D). The subsequent tracker pipeline consists of two main steps: 1. detection and classification of objects in the crops over time and 2. association of the detected objects over time to compute consistent identities. Detection and classification is done by the YOLOv7 framework, which has been used successfully on a wide range of domains including tiny insect detection [29] and medical imaging [30] and requires only moderate computing capabilities. For our experiments in Sec. 2 we used the baseline YOLOv7 model pretrained on the COCO dataset [31] and finetuned it on our custom Hatching-Box dataset (see Sec. 2.2). Loading and inference of our model is performed using the ONNX [32] framework to make use of platform-dependent acceleration methods. This decreases the inference time and provides an interface which makes it possible to also load other models to use in HB-server if desired. To further speed up the inference and the tracking process off-the-shelf GPU computing hardware is recommended.

### v3

@@ 4.2.4 Tracking
To analyse the images recorded by the HB-client we implemented an adapted tracking algorithm, which is accessible in the HB-server. This algorithm firstly separates the image into individual rearing vial crops using a one-time calibration step, which automatically generates a mask for each vial (see Appendix E). The subsequent tracker pipeline consists of two main steps: 1. detection and classification of objects in the crops over time and 2. association of the detected objects over time to compute consistent identities. Detection and classification is done by the YOLOv7 framework, which has been used successfully on a wide range of domains including tiny insect detection [31] and medical imaging [32] and requires only moderate computing capabilities. For our experiments in Sec. 2 we used the baseline YOLOv7 model pretrained on the COCO dataset [33] and finetuned it on our custom Hatching-Box dataset (see Sec. 2.2). Loading and inference of our model is performed using the ONNX [34] framework to make use of platform-dependent acceleration methods. This decreases the inference time and provides an interface which makes it possible to also load other models to use in HB-server if desired. To further speed up the inference and the tracking process off-the-shelf GPU computing hardware is recommended.